I'm trying to solve a problem. I solved half of it, but I don't know how to continue.
This is the problem: "If the length of a rectangle is 1 foot less than twice its width and its perimeter is 34 feet, what is the length of the rectangle in feet?

Thank you for your help
Lisa

Answer by rothauserc(4718)

let x be the width(w), then length(l) is 2x-1
Perimeter(P) of a rectangle = 2l + 2w
34 = 2(2x-1) +2x
34 = 4x-2 +2x
6x = 36
x = 6 and
l = 2*6 -1 = 11
the length of the rectangle is 11 feet